### v1

# SDGF: Fusing Static and Multi-Scale Dynamic Correlations for Multivariate Time Series Forecasting

Abstract
Inter-series correlations are crucial for accurate multivariate time series forecasting, yet these relationships often exhibit complex dynamics across different temporal scales. Existing methods are limited in modeling these multi-scale dependencies and struggle to capture their intricate and evolving nature. To address this challenge, this paper proposes a novel Static-Dynamic Graph Fusion network (SDGF), whose core lies in capturing multi-scale inter-series correlations through a dual-path graph structure learning approach. Specifically, the model utilizes a static graph based on prior knowledge to anchor long-term, stable dependencies, while concurrently employing Multi-level Wavelet Decomposition to extract multi-scale features for constructing an adaptively learned dynamic graph to capture associations at different scales. We design an attention-gated module to fuse these two complementary sources of information intelligently, and a multi-kernel dilated convolutional network is then used to deepen the understanding of temporal patterns. Comprehensive experiments on multiple widely used real-world benchmark datasets demonstrate the effectiveness of our proposed model. Code is available at https://github.com/shaoxun6033/SDGFNet.
Index Terms—  Time series forecasting, Wavelet decomposition, Graph neural network
1 Introduction
Time Series Forecasting (TSF) is a crucial research topic in the field of machine learning. It demonstrates significant application value across numerous critical domains, including financial market analysis [1], smart energy management [2], meteorological disaster early warning [3], and intelligent transportation planning [4]. Accurate forecasting capability forms the cornerstone for scientific decision-making, optimal resource allocation, and proactive risk management in modern society.
In multivariate time series forecasting, it is equally important to accurately capture the temporal dynamics within individual sequences (intra-series dependencies) and to effectively model the interactions among different sequences (inter-series dependencies) [5]. In recent years, advanced architectures—from RNNs [6], LSTM [7], TimeMixer [8], and DLinear [9] to Transformer-based models such as Informer [10], Autoformer [11], and PatchTST [12]—have demonstrated remarkable success in capturing long-range temporal dependencies and intra-series patterns. Meanwhile, Graph Neural Networks (GNNs) [13] have shown significant advantages in handling graph-structured data due to their permutation invariance, local connectivity, and compositionality. By viewing multivariate time series from a graph perspective, GNNs provide a natural and effective framework for modeling complex inter-series dependencies. However, existing GNN-based methods exhibit notable limitations: they either heavily rely on a predefined, prior-knowledge-based graph [14], which is unrealistic in many general scenarios, or—even when adopting adaptive graph learning—typically produce a single global static graph [15]. This “one-size-fits-all” approach overlooks the crucial fact that inter-variable correlations may exhibit markedly different patterns across multiple temporal scales. Although some efforts have attempted to address this issue [5, 16], they still struggle to fully capture the complex and dynamically evolving inter-series dependencies.
To address these challenges, we introduce Multi-level Wavelet Decomposition (MWD) to decompose the original multivariate time series into sub-sequences at different frequencies and temporal scales, thereby capturing multi-granularity inter-variable relationships. Based on this, we propose the Static-Dynamic Graph Fusion (SDGF) Network, which models long-term stable dependencies via a global static graph and learns dedicated dynamic graphs for each wavelet scale to capture scale-dependent relations.
Our main contributions are: (1)We propose the SDGF Network, which simultaneously models static and dynamic scale-dependent inter-series relations. (2)We introduce an attention-gated fusion mechanism to effectively integrate multi-scale dynamic graph convolutions with a global static graph. (3)We validate the effectiveness of our model on multiple benchmark datasets, showing superior predictive performance over existing state-of-the-art methods.
2 RELATED WORK
Deep learning-based time series forecasting has become the mainstream in research and applications, with MLPs and Transformers serving as the two dominant frameworks. MLP-based methods are favored for their simplicity and interpretability, such as TimeMixer [8], DLinear [9], and TimesNet [17]. Transformer-based methods, exemplified by Informer [10], Autoformer [11], and TimeXer [18], exploit self-attention to capture sequence dependencies, often combined with decomposition or Fourier transforms to enhance temporal modeling.
GNNs are specifically designed for graph-structured data, making them well-suited for modeling complex interactions and learning local patterns in time series. They were initially applied to spatio-temporal traffic forecasting [15] and skeleton-based action recognition [19]. Later, MTGNN [16] explored a learnable graph structure, while MSGNet [5] further advanced this direction by integrating frequency-domain analysis with adaptive graph convolutions, effectively capturing complex dependencies in multivariate time series data. However, these approaches typically rely on a limited number of graph structures or scales, and fail to incorporate global static structural relations.
3 METHOD
3.1 Problem Statement
Given a historical multivariate time series $X_{in}\in\mathbb{R}^{L\times N}$, where $L$ denotes the length of the historical window and $N$ is the number of variables, the task is to predict the future sequence $\hat{X}_{out}\in\mathbb{R}^{T\times N}$ for the next $T$ steps. To capture inter-variable dependencies, we employ wavelet decomposition to project the input series into multiple scales, and learn graph structures at each scale, yielding a set of adjacency matrices $\{A_{1},A_{2},\dots,A_{k}\}$ with $A_{j}\in\mathbb{R}^{N\times N}$. Finally, the forecasting function integrates the input sequence with the multi-scale graph structures to produce predictions as $\hat{X}_{out}=f(X_{in},\{A_{j}\}_{j=1}^{k})$.
3.2 Framework
Our proposed SDGF Network aims to capture complex inter-series correlations at different temporal scales through graph structure learning and enhance predictive accuracy by fusing dynamic and static graph structures. The overall architecture is illustrated in Fig. 1, and it consists of three core modules: (1) Graph Structure Learning: The input sequences are first normalized using RevIN normalization [20] to eliminate scale differences across series. Subsequently, Multi-level Wavelet Decomposition (MWD) is applied to extract multi-scale features, providing diverse frequency components for graph structure learning. The static graph structure (Priori Graph Structure) leverages prior knowledge or statistical correlations(Pearson Correlation Coefficient, PCC) to construct a fixed inter-series relationship graph, capturing long-term stable dependencies. In contrast, the dynamic graph structure adaptively learns variable inter-series relationships based on multi-scale features, capturing short-term or evolving dependencies. (2) Attention Gated Fusion: To effectively integrate static and dynamic graph convolution outputs, we employ an attention gated fusion module, which adaptively combines these features, yielding more accurate and flexible inter-series representations. (3) Temporal Feature Learning: The fused inter-series features are processed through multi-kernel dilated convolutions to extract high-level temporal representations, capturing both short- and long-term dependencies across multiple time scales. Finally, the learned features are fed into a lightweight MLP-based Output layer for nonlinear mapping to produce the final forecasts. In summary, SDGF Network leverages multi-scale graph structure learning, attention-based fusion, and advanced temporal feature extraction to provide an efficient and scalable framework for multivariate time series forecasting.
Fig. 1: Overall architecture of SDGF Network.
3.3 Graph Structure Learning
To eliminate distribution shift and scale differences, we first apply Reversible Instance Normalization (RevIN) to the input sequences, $\mathbf{X}_{\text{norm}}=\text{RevIN}(\mathbf{X},\text{mode}=\text{'norm'})$, where RevIN normalizes each instance by subtracting its mean and dividing by its standard deviation. Learnable affine parameters $(\gamma,\beta)$ ensure reversibility, allowing the predicted outputs to be mapped back to the original scale during inference.
Static Graph Structure (Prior Graph): The static graph captures long-term stable inter-series dependencies. Given the normalized input $\mathbf{X}_{\text{norm}}\in\mathbb{R}^{B\times L\times N}$, where $B$ is the batch size, $L$ is the sequence length, and $N$ is the number of variables, the adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ is computed based on the Pearson Correlation Coefficient (PCC):
$\mathbf{A}_{nm}=\frac{\text{Cov}(\mathbf{X}_{\text{norm},:,n},\mathbf{X}_{\text{norm},:,m})}{\sigma_{n}\sigma_{m}+\epsilon}$ (1)
where $\text{Cov}(\cdot,\cdot)$ is the covariance, $\sigma_{n}$ and $\sigma_{m}$ are the standard deviations of sequences $n$ and $m$, and $\epsilon=10^{-9}$ prevents division by zero. Averaging over the batch dimension and applying ReLU followed by softmax ensures non-negativity and normalization:
$\mathbf{A}=\text{Softmax}\big{(}\text{ReLU}(\text{Mean}_{B}(\mathbf{A}))\big{)}$ (2)
The static graph convolution propagation is formulated as:
$\mathbf{H}^{static}=\sum_{k=0}^{K}\Theta_{k}\Big{(}\alpha\mathbf{H}+(1-\alpha)\mathbf{A}\mathbf{H}\Big{)}$ (3)
where $\alpha\in(0,1)$ controls the residual contribution, $K$ is the propagation depth, and $\Theta_{k}$ are learnable parameters.
Dynamic Graph Structure (Wavelet Decomposition): To capture short-term or evolving dependencies, we perform multi-level wavelet decomposition (MWD) on the normalized input, i.e., $\mathbf{X}_{\text{norm}}=\sum_{l=1}^{L_{d}}\mathbf{X}^{(l)},\;\mathbf{X}^{(l)}\in\mathbb{R}^{B\times L_{l}\times N}$, where $L_{d}$ denotes the decomposition level.
For each sub-sequence $\mathbf{X}^{(l)}$, a dynamic adjacency matrix is generated and graph convolution is applied:
$\mathbf{A}^{dyn,(l)}=\text{Softmax}\Big{(}\tanh(W_{1}\mathbf{X}^{(l)})\cdot\tanh(W_{2}\mathbf{X}^{(l)})\Big{)}$ (4)
$\mathbf{H}^{dyn,(l)}=\sum_{k=0}^{K}\Theta^{(l)}_{k}\Big{(}\alpha\mathbf{H}^{(l)}+(1-\alpha)\mathbf{A}^{dyn,(l)}\mathbf{H}^{(l)}\Big{)}$ (5)
The multi-scale dynamic graph feature set is denoted as:
$\mathcal{H}^{dyn}=\{\mathbf{H}^{dyn,(1)},\mathbf{H}^{dyn,(2)},\dots,\mathbf{H}^{dyn,(L_{d})}\}$ (6)
3.4 Attention Gated Fusion
To effectively integrate static and dynamic graph features, we employ an Attention Gated Fusion module. The input consists of the static graph representation $\mathbf{H}^{static}\in\mathbb{R}^{B\times D\times N}$ and the multi-scale dynamic graph set $\mathcal{H}^{dyn}$.
All graph representations are stacked:
$\mathcal{H}=[\mathbf{H}^{static},\mathbf{H}^{dyn,(1)},\dots,\mathbf{H}^{dyn,(L_{d})}]\in\mathbb{R}^{B\times(L_{d}+1)\times D\times N}$ (7)
Each graph is pooled over nodes:
$h_{i}=\text{AvgPool}_{N}(\mathcal{H}_{i}),\quad h_{i}\in\mathbb{R}^{B\times D}$ (8)
Attention weights are computed using a learnable query vector $q\in\mathbb{R}^{D}$:
$\alpha_{i}=\frac{\exp(q^{\top}W_{k}h_{i})}{\sum_{j}\exp(q^{\top}W_{k}h_{j})},\quad\mathbf{H}^{fusion}=\sum_{i=0}^{L_{d}}\alpha_{i}\mathcal{H}_{i}$ (9)
3.5 Temporal Feature Learning
The fused representation $\mathbf{H}^{fusion}\in\mathbb{R}^{B\times D\times N}$ is processed by a multi-kernel dilated convolution Inception module to capture multi-scale temporal dependencies and includes a residual connection. The computation is formulated as:
$\mathbf{H}_{concat}=\big{[}\text{Conv1D}_{k,d}(\mathbf{H}^{fusion})\big{]}_{k\in\{3,5\},d\in\{1,2\}}$ (10)
$\mathbf{H}_{out}=\text{Conv1D}_{1}(\mathbf{H}_{concat})+\text{Conv1D}_{1}(\mathbf{H}^{fusion})$ (11)
$\mathbf{H}^{temp}=\text{LayerNorm}(\mathbf{H}_{out})$ (12)
This module captures both short-term local patterns and long-term temporal dependencies, providing high-level temporal features for the subsequent prediction module output layer.
Table 1: Quantitative evaluation of time series forecasting models with $L=96$ and prediction length $T\in\{96,192,336,720\}$.
Models Ours TimesNet DLinear MSGNet MTGNN Autoformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather
96 0.169 0.215 0.172 0.220 0.196 0.255 0.163 0.212 0.171 0.231 0.266 0.336 0.184 0.227
192 0.211 0.256 0.219 0.261 0.237 0.296 0.212 0.254 0.215 0.274 0.307 0.367 0.234 0.265
336 0.269 0.297 0.280 0.306 0.283 0.335 0.272 0.299 0.266 0.313 0.359 0.395 0.284 0.301
720 0.346 0.348 0.365 0.359 0.345 0.381 0.350 0.348 0.344 0.375 0.419 0.428 0.356 0.349
ETT*
96 0.343 0.388 0.384 0.402 0.386 0.400 0.390 0.411 0.440 0.450 0.449 0.459 0.460 0.447
192 0.361 0.407 0.436 0.429 0.437 0.432 0.442 0.442 0.449 0.433 0.500 0.482 0.512 0.477
336 0.370 0.419 0.438 0.469 0.481 0.459 0.480 0.468 0.598 0.534 0.521 0.496 0.546 0.496
720 0.454 0.486 0.521 0.500 0.519 0.516 0.494 0.488 0.685 0.620 0.514 0.512 0.544 0.517
Electricity
96 0.179 0.268 0.162 0.272 0.197 0.282 0.165 0.274 0.211 0.305 0.201 0.317 0.190 0.296
192 0.181 0.270 0.184 0.289 0.196 0.285 0.184 0.292 0.225 0.319 0.222 0.334 0.199 0.304
336 0.198 0.293 0.198 0.300 0.209 0.301 0.195 0.302 0.247 0.340 0.231 0.338 0.217 0.319
720 0.248 0.325 0.220 0.320 0.245 0.333 0.231 0.332 0.287 0.373 0.254 0.361 0.258 0.352
Exchange
96 0.087 0.207 0.107 0.234 0.088 0.218 0.102 0.230 0.267 0.378 0.197 0.323 0.088 0.205
192 0.181 0.303 0.226 0.344 0.176 0.315 0.195 0.317 0.590 0.578 0.300 0.369 0.176 0.299
336 0.342 0.425 0.367 0.448 0.313 0.427 0.359 0.436 0.939 0.749 0.509 0.524 0.301 0.397
720 0.869 0.705 0.964 0.746 0.839 0.695 0.940 0.738 1.107 0.834 1.447 0.941 0.901 0.714
$1^{st}$ Count 18 3 1 4 1 0 3
4 EXPERIMENTS
4.1 Experimental Setup
Datasets. To evaluate the effectiveness of our proposed method, we conducted experiments on seven benchmark time series datasets: ETT (h1, h2, m1, m2), Exchange-Rate, Electricity and Weather. For each dataset, we split the data into training, validation, and test sets with a 7:2:1 ratio.
Baselines. We evaluate our method against seven SOTA models, including GNN-based MSGNet [5] and MTGNN [16], CNN-based TimesNet [17], Transformer-based PatchTST [12] and Autoformer [11], and the MLP-based DLinear [9].
Implementation. All experiments were implemented using PyTorch on a single NVIDIA GTX 4090 24GB GPU. The model was optimized with the Adam optimizer and $Batch=32$. Following standard time series forecasting protocols, the lookback length was set to $L=96$, and $T\in\{96,192,336,720\}$. Forecast performance was evaluated using Mean Squared Error (MSE) and Mean Absolute Error (MAE)  [5, 9, 11, 17].
4.2 Results and Analysis
From the main results in Table 1, our model achieves the best or second-best performance in most test tasks (best results in bold, second-best underlined), demonstrating its effectiveness and robustness. The advantage is particularly notable on the Weather and ETT* datasets, which exhibit pronounced multi-scale and periodic patterns. Specifically, on ETT*, our model attains the best MSE and MAE across all four prediction horizons. In comparison, GNN-based methods such as MTGNN and MSGNet perform worse, as they typically learn a single global static graph or fail to dynamically integrate graph structures with the inherent multi-scale characteristics of the signals, limiting their ability to capture evolving inter-variable dependencies. For visualization, partial prediction results are illustrated in Figure 2. Note that ETT* represents ETTh1, and other experimental results can be reproduced using the publicly available code.
(a) Electricity, output-96
(b) Electricity, output-192
(c) Electricity, output-336
(d) Electricity, output-720
Fig. 2: Visualization comparison of forecasting results.
4.3 Ablation Study
The ablation results on ETTh1, Exchange, and Weather (Table 2) demonstrate the importance of each component in SDGFNet. Removing graph structure learning (w/o-GSL) leads to a clear performance drop, highlighting the necessity of modeling both static and dynamic inter-series dependencies. Similarly, excluding the attention-gated fusion module (w/o-GF) degrades accuracy, verifying the effectiveness of integrating static and dynamic graphs. Moreover, removing temporal feature learning (w/o-TFL) also reduces performance, showing that multi-kernel dilated convolutions are essential for capturing cross-scale temporal dependencies. Overall, these results confirm that all three modules are indispensable to the superior performance of SDGFNet.
Table 2: Ablation analysis of SDGFNet.
Dataset ETTh1 Exchange Weather
Metric MSE MAE MSE MAE MSE MAE
SDGFNet 0.343 0.388 0.087 0.207 0.169 0.215
w/o-GSL 0.366 0.401 0.091 0.219 0.176 0.224
w/o-GF 0.371 0.397 0.102 0.221 0.185 0.241
w/o-TFL 0.368 0.409 0.099 0.215 0.181 0.237
5 CONCLUSION
In this paper, we introduced the SDGF Network, a novel dual-branch architecture for multivariate time series forecasting. By explicitly decoupling and fusing static global correlations with multi-scale dynamic dependencies learned via wavelet decomposition, our model captures a richer set of inter-series relationships. An attention mechanism intelligently integrates these complementary features. Extensive experiments on several benchmark datasets demonstrate the effectiveness and competitive performance of our proposed method.

### v2

@@
 Abstract
-Inter-series correlations are crucial for accurate multivariate time series forecasting, yet these relationships often exhibit complex dynamics across different temporal scales.
+Accurate multivariate time series forecasting hinges on inter-series correlations, which often evolve in complex ways across different temporal scales.
 Existing methods are limited in modeling these multi-scale dependencies and struggle to capture their intricate and evolving nature.
@@
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Weather
-96 0.169 0.215 0.172 0.220 0.196 0.255 0.163 0.212 0.171 0.231 0.266 0.336 0.184 0.227
+Weather 96 0.169 0.215 0.172 0.220 0.196 0.255 0.163 0.212 0.171 0.231 0.266 0.336 0.184 0.227
 192 0.211 0.256 0.219 0.261 0.237 0.296 0.212 0.254 0.215 0.274 0.307 0.367 0.234 0.265
@@
 720 0.346 0.348 0.365 0.359 0.345 0.381 0.350 0.348 0.344 0.375 0.419 0.428 0.356 0.349
-ETT*
-96 0.343 0.388 0.384 0.402 0.386 0.400 0.390 0.411 0.440 0.450 0.449 0.459 0.460 0.447
+ETT* 96 0.343 0.388 0.384 0.402 0.386 0.400 0.390 0.411 0.440 0.450 0.449 0.459 0.460 0.447
 192 0.361 0.407 0.436 0.429 0.437 0.432 0.442 0.442 0.449 0.433 0.500 0.482 0.512 0.477
@@
 720 0.454 0.486 0.521 0.500 0.519 0.516 0.494 0.488 0.685 0.620 0.514 0.512 0.544 0.517
-Electricity
-96 0.179 0.268 0.162 0.272 0.197 0.282 0.165 0.274 0.211 0.305 0.201 0.317 0.190 0.296
+Electricity 96 0.179 0.268 0.162 0.272 0.197 0.282 0.165 0.274 0.211 0.305 0.201 0.317 0.190 0.296
 192 0.181 0.270 0.184 0.289 0.196 0.285 0.184 0.292 0.225 0.319 0.222 0.334 0.199 0.304
@@
 720 0.248 0.325 0.220 0.320 0.245 0.333 0.231 0.332 0.287 0.373 0.254 0.361 0.258 0.352
-Exchange
-96 0.087 0.207 0.107 0.234 0.088 0.218 0.102 0.230 0.267 0.378 0.197 0.323 0.088 0.205
+Exchange 96 0.087 0.207 0.107 0.234 0.088 0.218 0.102 0.230 0.267 0.378 0.197 0.323 0.088 0.205
 192 0.181 0.303 0.226 0.344 0.176 0.315 0.195 0.317 0.590 0.578 0.300 0.369 0.176 0.299
@@
 720 0.869 0.705 0.964 0.746 0.839 0.695 0.940 0.738 1.107 0.834 1.447 0.941 0.901 0.714
-$1^{st}$ Count 18 3 1 4 1 0 3
+$1^{st}$ Count 18 3 4 6 0 0 5
 4 EXPERIMENTS